OpenStudy (anonymous):

What is the sum of the finite arithmetic series? (–5) + 0 + 5 + 10 + ... + 65

OpenStudy (lacypennelll):

SO is that the full prob? (-5) + 0 + 5 + 10 + ... + 65?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

HELP ANYONE ?

OpenStudy (anonymous):

@phi Please help me

OpenStudy (anonymous):

[n/2(2A1+(n-1)d]

OpenStudy (anonymous):

What does that mean ? Does that go to the problem to get my answer ?

OpenStudy (anonymous):

first get the common difference of the sequence @Kayy_Drizzyy

OpenStudy (anonymous):

yes thats the formula to get the series.

OpenStudy (anonymous):

Oh kay. Hold on one second. let me figure it out.

OpenStudy (anonymous):

Assuming that -5 is a_0, we can characterize the sequence in the following way:\[a_n = (n-1) \cdot 5\] Hence, to find out the index of the last term in the sum, we solve the following equation:\[65 = (i - 1) \cdot 5\]\[\frac{65}{5} = i - 1\]\[i = \frac{65}{5} + 1 = 14\] So the sum above equals \[\sum_{n=0}^{14}a_n = \sum_{n=0}^{14}(n-1)\cdot 5 = 5\cdot \sum_{n=0}^{14}(n-1) = 5\cdot (\sum_{n=0}^{14}n - \sum_{n=0}^{14}1)\]\[= 5\cdot \sum_{n=0}^{14}n - 5\cdot 14 = 5\cdot \frac{14\cdot 15}{2} - 5 \cdot 14 = 455\]

OpenStudy (anonymous):

Thank you @Stiwan that really helps a lot.

OpenStudy (anonymous):

\[18Sigmat=1\]

OpenStudy (anonymous):

\[\sum_{t=1}^{18}\]

OpenStudy (anonymous):

I'm sorry, I really don't get what that's supposed to mean

OpenStudy (anonymous):

i was testing something out sorry

OpenStudy (anonymous):

is this problem is done?

OpenStudy (anonymous):

yes